Circular No. 5852

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

RX J1940.2–1025 J. Patterson and J. Halpern, Columbia University, write: “Photometry

by D. R. Skillman shows variations from V = 16.9 to 16.0, with a strict period of 12 150 s, equal within the 0.1-percent measurement error to the x-ray period previously attributed to NGC 6814. The measurements were made over a 12-day baseline with the 0.32-m CBA reflector (Skillman 1993, Sky Tel. 85, No. 5, 83), and maximum light followed the ephemeris HJD 2449199.608 + 0.14062E.”

Corrigendum. On IAUC 5850, line 7, for count rate of 1.4 counts/s. read count rate of 0.2 counts/s.

V HYDRAE T. L. Evans, South African Astronomical Observatory, reports: “V

Hya is a semiregular variable, of spectral type N, with a period of 531 days superimposed on a large amplitude variation with a period of 17-18 yr. Spectroscopic exposure times indicate a decline of 2 mag in visible light between April and July, which may indicate the start of the deep minimum that was expected about 1994. Infrared photometry by B. S. Carter on Aug. 8 gives J = 4.0, H = 2.0, K = 0.5, and L = –1.3. This is 2 mag fainter in J and 1.1 mag redder in J–K than in mid-1975 but agrees well with an observation in early 1978, when V Hya would have been in its last deep minimum. Spectroscopic monitoring every few months since 1986 December shows that spectroscopic changes started in 1992 July: NaD absorption weakened and became net emission in 1992 December and 1993 April, reverting to absorption with emission cores in 1993 July. The C_{2} emission at 473.7 nm strengthened considerably in this period and strong emission appeared for the first time in the 516.5-nm bandhead of C_{2}. Hα emission showed its usual variation in the 531-day period throughout. This suggests that V Hya ejected gas, which then shone in NaD and C_{2} by resonance emission from the middle of 1992, and that dust condensation a year later has caused the fading of the star. The jet and accretion disk have also faded, ruling out an explanation in terms of an eclipse by dust associated with the latter. Observations in the far-infrared and microwave regions of the spectrum are needed to search for other manifestations of mass-loss.”

1993 September 2 Daniel W. E. Green